OpenStudy (mindblast3r):

1/3x = 12 solve for x

OpenStudy (freckles):

you want to isolate x but oh no x is being multiplied by 1/3 to undo that multiplication you can divide both sides by 1/3

OpenStudy (mathstudent55):

Do you mean \(\dfrac{1}{3} x = 12\) ?

OpenStudy (mindblast3r):

yes @mathstudent55

OpenStudy (godlovesme):

 can u solve from there?

OpenStudy (mindblast3r):

36

OpenStudy (mindblast3r):

can't i just add 2/3 to both sides?

OpenStudy (mathstudent55):

Ok. Since x is being multiplied by 1/3 and you want x alone, multiply both sides by the reciprocal of 1/3, which is 3.

OpenStudy (godlovesme):

great :D

OpenStudy (mathstudent55):

No. Adding 2/3 will not help.

OpenStudy (godlovesme):

1/3 is multiplied so u gotta multiply it by the reciprocal

OpenStudy (mathstudent55):

\(\dfrac{1}{3} x = 12\) Add 2/3 to both sides: \(\dfrac{1}{3} x + \dfrac{2}{3} = 12 + \dfrac{2}{3}\) \(\dfrac{1}{3} x + \dfrac{2}{3} = 12 \dfrac{2}{3}\) Since 1/3 x and 2/3 are not like terms, they can't be combined together, so you have just complicated the equation instead of solving it.

OpenStudy (mindblast3r):

hmm, i'm thinking

OpenStudy (godlovesme):

where did u get 2/3 from @mindblast3r

OpenStudy (mathstudent55):

The way to solve an equation is to do the opposite operation of what is being done to x. The opposite of multiplying by 1/3 is divide by 1/3, but dividing by 1/3 is the same as multiplying by 3.

OpenStudy (mathstudent55):

I assume the idea was to do 1/3 + 2/3 = 1

OpenStudy (mindblast3r):

yes lol @mathstudent55

OpenStudy (mindblast3r):

well i know adding 2/3 was wrong but, my main intentions of the question was to find out why it's wrong, i just didn't know how to phrase the question.

OpenStudy (mathstudent55):

I hope you understand it now.

OpenStudy (mindblast3r):

The way to solve an equation is to do the opposite operation of what is being done to x. is this like an official rule?

OpenStudy (mathstudent55):

No, I never heard of it as an officvial rule, but that's how we solve equations.

OpenStudy (mindblast3r):

who is "we", humans?

OpenStudy (mathstudent55):

Yes, "we" are people. Here is something for youto think about. Take an equation such as 2x + 5 = 15

OpenStudy (godlovesme):

i'll let @mathstudent55 help u :D

OpenStudy (mathstudent55):

You want x alone, but there is a 5 being added and a 2 being multiplied.

OpenStudy (mindblast3r):

thank you veryyyyyyyyyyyyyy much guys i <3 all of you

OpenStudy (mathstudent55):

How do you figure out which exact operations to do to get rid of the plus 5 and the times 2?

OpenStudy (mathstudent55):

Think of just the left side, 2x + 5. If you were asked to evaluate the left side for a number, such as x = 10, what would you do?

OpenStudy (mindblast3r):

1. subtract 5 from both sides 2. divide both sides by 2 3. x = 5

OpenStudy (mathstudent55):

First, you replace x with 10. 2x + 5 2(10) + 5 Now you do the multipluication first, 2 * 10 20 + 5 Finally youi do the aqddition, 20 + 5 25

OpenStudy (mathstudent55):

You are correct.

OpenStudy (mathstudent55):

Now look at what I did to evaluate the left side. 1) I first multiplied by 2 2) then I added 5

OpenStudy (mathstudent55):

To solve the equation, you do the opposite operation in the opposite order. To Evaluate: To Solve: multiply by 2 subtract 5 add 5 divide by 2

OpenStudy (mindblast3r):

okay i understand now. thank you :D